# SECURE: Benchmarking Large Language Models for Cybersecurity Advisory

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated potential in cybersecurity applications but have also caused lower confidence due to problems like hallucinations and a lack of truthfulness. Existing benchmarks provide general evaluations but do not sufficiently address the practical and applied aspects of LLM performance in cybersecurity-specific tasks. To address this gap, we introduce the SECURE (Security Extraction, Understanding & Reasoning Evaluation), a benchmark designed to assess LLMs performance in realistic cybersecurity scenarios. SECURE includes six datasets focussed on the Industrial Control System sector to evaluate knowledge extraction, understanding, and reasoning based on industry-standard sources. Our study evaluates seven state-of-the-art models on these tasks, providing insights into their strengths and weaknesses in cybersecurity contexts. We also offer recommendations for improving LLMs reliability as cyber advisory tools and release our benchmark datasets and framework for community use at https://github.com/aiforsec/SECURE.
@@
Large Language Models, Security, Cyber Advisory, Dataset.
@@ I Introduction
    Assessment of cybersecurity knowledge: We evaluate the cybersecurity knowledge of LLMs in assisting security analysts. To achieve this, we develop a series of benchmark tasks specifically for evaluating LLMs in the context of ICS cyber advisory (see Section III). These tasks are based on comprehensive protocol specifications detailing communication standards and procedures, historical cyber threats and attacks, CVEs with contextual information on exploitability and severity, and detailed remediation strategies, including patch notes and mitigation steps from MITRE.
@@
    Evaluation of LLMs: Recent works have shown that customized LLMs introduce a loss of safety measures [16], & increased hallucinations [17]. In addition, the computational & technical requirements of building customized LLMs can push researchers & industries to use commercial LLMs. However, before we can confidently claim the application of general-purpose LLMs in a specific domain, we first need to evaluate it on domain-specific tasks. For this, we first evaluate 7 open and closed source state-of-the-art models (ChatGPT-4 [1], ChatGPT-3.5 [18], Llama3-70b [19], Llama3-8b [20], Gemini-Pro [21], Mistral-7B [22] , Mixtral-8x7b [23]) on our benchmark tasks (see Section IV). We then customize Llama3-8b using Retrieval Augmented Generation (RAG) framework and fine-tuning and evaluate the LLMs on our benchmark (see Section VI).
@@
    Insights and Recommendations: Our findings demonstrate that while LLMs demonstrate some capability in cybersecurity tasks, their use as advisory tools requires careful consideration. We provide insightful observations and propose recommendations to enhance their usability (see Section V).
@@
    Benchmark Dataset and Framework: We provide our benchmark datasets (MAET: Mitre Attack Extraction Task, CWET: Common Weakness Extraction Task, KCV: Knowledge test on Common Vulnerabilities, VOOD: Vulnerability Out-of-Distribution, RERT: Risk Evaluation Reasoning Task, and CPST: CVSS Problem Solving Task) and framework to the security community so that they can be openly used to evaluate future LLMs in various security applications.
@@ III-A Modeling
We emphasize knowledge modeling in designing our evaluation benchmark. Knowledge, which includes both facts and skills, is a core indicator of intelligence [43]. Prior research shows that knowledge-intensive tasks can reliably evaluate the capabilities of LLMs [9, 44]. For cyber-advisory LLM, we aim to assess not only its ability to retrieve known facts but also gauge its proficiency in a) using context to answer questions and b) performing reasoning based on given some knowledge source. To create a robust evaluation benchmark, we focus on the following critical abilities of LLMs in handling knowledge-intensive tasks:
Extraction: Knowledge extraction tasks are crucial for measuring the ability of a language model to access its vast knowledge base and accurately recall specific facts [44]. As a cybersecurity advisor, an LLM could be asked to retrieve information on various security architectures, incidents, best practices, and historical data on known vulnerabilities. Accurate extraction ensures the LLM can provide reliable and precise information quickly, which is essential for addressing security concerns, preventing breaches, and supporting security professionals in their decision-making processes.
@@ III-B1 Extraction Task
Prompt A: From the following URL ’\${URL}’, generate a set of MCQs (zero to five) for ‘novices’ and similarly for ‘experts’ with four possible answers each. … Return the output in CSV format (tab separated) for the responses with the following nine columns: URL, Level (Novice or Expert), Question, Option A, Option B, Option C, Option D, Correct Answer (A, B, C or D), Explanation…
@@ III-B2 Understanding Task
Our second task dataset is designed to evaluate the ability of LLM to comprehend and understand the security-related text. Given the continuously evolving landscape of cybersecurity, it is crucial that LLMs can assimilate new information and use it to generate accurate responses. For this purpose, we utilize the CVEs (Common Vulnerabilities and Exposures) published in 2024, available at CVE project repository [13]. CVE is better for this task since they are always updated but CWEs are more appropriate for the extraction task as they are updated less frequently.
@@
For this task, we generated 466 boolean questions. This dataset serves as a robust benchmark for evaluating the comprehension capabilities of LLMs in cybersecurity, mainly focusing on their ability to adapt to and reason about newly encountered vulnerabilities.
@@ IV Experiments & Results
We evaluate 7 state-of-the-art LLMs varying in parameter size, organization, and access (See Appendix -A and Table VIII for detail and comparative summary). We pick both open-source and closed models based on LLM leaderboards (ChatArena [55]). Open-source models allow the download of full model weights whereas closed models provide an API for restricted access. We evaluate our benchmark against the following models:
@@ IV-B Evaluation Metrics
We employ a range of evaluation metrics tailored to the specific nature of each task within our benchmarking framework: accuracy for MAET, CWET, VOOD and KCV, ROGUE-L [54] for RERT and mean absolute deviation (MAD) for CPST. Because of the page-limit, we explain the metrics in detail in Appendix -B.
@@ VI-B Fine-tuning
Fine-tuning is the process of taking a pretrained model and training it further on a specific dataset to adapt it to particular tasks. The goal of this ‘tuning’ is to help the model learn the nuances of the new data. We follow the following steps:
@@ VIII Conclusion
LLMs are rapidly transforming the landscape of AI, with vast potential applications in cybersecurity. However, concerns regarding their reliability, and understanding, particularly within the sensitive domain of cybersecurity, remain paramount. The SECURE benchmark, introduced in this paper, aims to address these concerns by providing a holistic framework for evaluating the capabilities of LLMs in a cybersecurity advisory context. Our experiments and analysis demonstrate the potential of LLMs as valuable tools for cybersecurity professionals. However, our findings also underscore the need for caution, especially when handling reasoning tasks. By open-sourcing our benchmark datasets, we invite the research community to contribute to their improvement, enhancing LLMs reliability and paving the way for responsible and beneficial large language models in cybersecurity.
Acknowledgement
This work is supported by Grant H98230-21-1-0317.

### v4

# Rochester, USA

@@ Abstract
Large Language Models (LLMs) have demonstrated potential in cybersecurity applications but have also caused lower confidence due to problems like hallucinations and a lack of truthfulness. Existing benchmarks provide general evaluations but do not sufficiently address the practical and applied aspects of LLM performance in cybersecurity-specific tasks. To address this gap, we introduce the SECURE (Security Extraction, Understanding & Reasoning Evaluation), a benchmark designed to assess LLMs performance in realistic cybersecurity scenarios. SECURE includes six datasets focused on the Industrial Control System sector to evaluate knowledge extraction, understanding, and reasoning based on industry-standard sources. Our study evaluates seven state-of-the-art models on these tasks, providing insights into their strengths and weaknesses in cybersecurity contexts. We also offer recommendations for improving LLMs reliability as cyber advisory tools and release our benchmark datasets and framework for community use at https://github.com/aiforsec/SECURE.
@@
Large Language Models, Cybersecurity, Benchmarking, Dataset, Industrial Control System.
@@ I Introduction
    Assessment of Cybersecurity Knowledge: We evaluate the cybersecurity knowledge of LLMs in terms of their capacity to assist security analysts. Specifically, we design a set of benchmark tasks tailored for the evaluation of LLMs within the context of Industrial Control Systems (ICS) cyber advisories (see Section III). These tasks are grounded in comprehensive protocol specifications, which include communication standards, Common Vulnerabilities and Exposures (CVEs) with contextual details on exploitability and severity, and remediation strategies, such as patch notes and mitigation steps informed by the MITRE framework.
@@
    Evaluation of LLMs: Recent works have shown that customized LLMs introduce a loss of safety measures [16], & increased hallucinations [17]. In addition, the computational and technical requirements of building customized LLMs can push researchers and industries to use commercial LLMs. However, before we can confidently claim the application of general-purpose LLMs in a specific domain, we first need to evaluate it on domain-specific tasks. To this end, we assess the performance of seven state-of-the-art, open-source, and proprietary LLMs on our cybersecurity benchmark tasks (see Section IV). These models include ChatGPT-4 [1], ChatGPT-3.5 [18], Llama3-70B [19], Llama3-8B [20], Gemini-Pro [21], Mistral-7B [22], and Mixtral-8x7B [23]. Following this, we customize Llama3-8B using the Retrieval-Augmented Generation (RAG) framework and fine-tuning and re-evaluate its performance on the benchmark (see Section VI).
@@
    Insights and Recommendations: Our evaluation reveals that, although LLMs exhibit some competence in cybersecurity tasks, their application as advisory tools demands careful consideration. We present key insights and offer recommendations to improve their practical usability (see Section V).
@@
    Benchmark Dataset and Framework: We present a comprehensive set of benchmark datasets—MAET (Mitre Attack Extraction Task), CWET (Common Weakness Extraction Task), KCV (Knowledge Test on Common Vulnerabilities), VOOD (Vulnerability Out-of-Distribution), RERT (Risk Evaluation Reasoning Task), and CPST (CVSS Problem Solving Task)—along with an evaluation framework. These resources are made available to the security community to facilitate the open evaluation of future LLMs across various security-related tasks.
@@ III-A Modeling
We emphasize knowledge modeling in designing our evaluation benchmark. Knowledge, which includes both facts and skills, is a core indicator of intelligence [43]. Prior research shows that knowledge-intensive tasks can reliably evaluate the capabilities of LLMs [9, 44]. For cyber-advisory LLM, we aim to assess not only its ability to retrieve known facts but also gauge its proficiency in a) using context to answer questions and b) performing reasoning based on given knowledge source. To create a robust evaluation benchmark, we focus on the following critical abilities of LLMs in handling knowledge-intensive tasks:
Extraction: Knowledge extraction tasks are critical for assessing a language model’s ability to access and accurately retrieve specific information from its extensive knowledge base [44]. In the context of cybersecurity advisory, an LLM may be required to provide information on various security frameworks, incidents, best practices, and historical data on known vulnerabilities. Accurate extraction is essential to ensure the LLM can deliver reliable and precise information promptly, which is crucial for addressing security concerns, preventing breaches, and assisting security professionals in their decision-making processes.
@@ III-B1 Extraction Task
Prompt A: From the following URL ‘\${URL}’, generate a set of MCQs (zero to five) for ‘novices’ and similarly for ‘experts’ with four possible answers each. … Return the output in CSV format (tab separated) for the responses with the following nine columns: URL, Level (Novice or Expert), Question, Option A, Option B, Option C, Option D, Correct Answer (A, B, C or D), Explanation…
@@ III-B2 Understanding Task
Our second task is designed to evaluate the ability of LLM to comprehend and understand the security-related text. Given the continuously evolving landscape of cybersecurity, it is crucial that LLMs can assimilate new information and use it to generate accurate responses. For this purpose, we utilize the CVEs (Common Vulnerabilities and Exposures) published in 2024, available at CVE project repository [13]. CVE is better for this task since they are always updated but CWEs are more appropriate for the extraction task as they are updated less frequently.
@@
For both KCV and VOOD, we generated 466 boolean questions. These datasets serve as a robust benchmark for evaluating the comprehension capabilities of LLMs in cybersecurity, mainly focusing on their ability to adapt to and reason about newly encountered vulnerabilities.
@@ IV Experiments & Results
We evaluate 7 state-of-the-art LLMs varying in parameter size, organization, and access (See Appendix -A and Table VIII for more detail). We pick both open-source and closed models based on LLM leaderboards (ChatArena [55]). Open-source models allow the download of full model weights whereas closed models provide an API for restricted access. We evaluate our benchmark against the following models:
@@ IV-B Evaluation Metrics
We employ a range of evaluation metrics tailored to the specific nature of each task within our benchmarking framework: accuracy for MAET, CWET, VOOD and KCV, ROGUE-L [54] for RERT and mean absolute deviation (MAD) for CPST. We explain the metrics in detail in Appendix -B.
@@ VI-B Fine-tuning
Fine-tuning is the process of taking a pretrained model and training it further on a specific dataset to adapt it to particular tasks [62]. The goal of this ‘tuning’ is to help the model learn the nuances of the new data. We follow the following steps:
@@ VIII Conclusion
LLMs are rapidly transforming the landscape of artificial intelligence, offering vast potential applications in cybersecurity. However, concerns surrounding their reliability and understanding remain paramount, particularly within the sensitive and high-stakes cybersecurity domain. The SECURE benchmark, introduced in this paper, addresses these concerns by providing a comprehensive framework for assessing the capabilities of LLMs in a cybersecurity advisory context. Our benchmark encompasses diverse tasks, evaluating knowledge extraction and reasoning abilities to ensure a thorough assessment. Through our experiments and analysis, we demonstrate the potential of LLMs as valuable tools for cybersecurity professionals. However, our findings also highlight the need for caution, especially in handling complex reasoning tasks. By open-sourcing our benchmark datasets, we invite the broader research community to contribute to their refinement, thereby enhancing the reliability of LLMs and paving the way for their responsible and effective use in cybersecurity applications.